A plane flies 300 km due north, 400 km due east, and then 500 km due south. How far is the plane from its starting point? State to the nearest km. My answer is approximately 860 km. Please confirm if this is correct.
Using Distance Formula: D2 = (x2-x1)+(y2-y1)
Starting Point: 0
East: positive x
West: negative x
North: positive y
South: negative y
Start: (0,0)
300 N: (0,0)+(300,0)=(300,0)
400 E: (300,0)+(400,0)=(700,0)
500 S: (700,0)+(0,-500)=(700,-500)
The plane is 700km east and -500km south of starting point.
D2= (700-0)2+ (0-500)2
D2= (700)2 + (-500)2
D2= 490,000 + 250,000
D2= 740,000
D = 860km
The plane is approximately 860km from its starting point.
Thank you, in advance, for your assistance.

Answer by mananth(16949)

The plane has taken a rectangular path
300 North
Turn right to east 400
then down south 500
it is 200 miles more south than the starting point
the horizontal distance is 400 km
Apply Pythagoras theorem
400^2+200^2= distance ^2
160000+40000=d^2
200000= d^2
d= sqrt (400000)
d= 447.2 km from starting point